[Brahms-l] Shift report 20040310 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Wed Mar 10 2004 - 08:00:41 EST
------------------------------------------------------------
Date: Wed Mar 10 08:00:40 2004, Shift: 00:00-08:00
Supervisor: CEJ, with: FH
Log book#15 pages 81 - 81

Shift Summary:
------------------------------------------------------------
1) One hour of beam from 3:30 to 4:45 (dumped because 
   of blue QLI. Beam again at 7:00

2) Pedestal runs 10872 and 10873 (TPC)
   Data runs: 
   10870, tr2: 65K, tr3: 72K
   10871, tr2: 29K, tr3: 43K
   10874, tr2: 50K, tr3: 45K
   10875, still running...

3) T3-2-X1 distribution looks bad. No other problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Mar 10 08:01:07 2004

This archive was generated by hypermail 2.1.8 : Wed Mar 10 2004 - 08:01:23 EST